Average Rate of Return Method, Net Present Value Method, and Analysis for a service company

The capital investment committee of Arches Landscaping Company is considering two capital investments. The estimated income from operations and net cash flows from each investment are as follows:

Front-End Loader Greenhouse
Year Income from
Operations
Net Cash
Flow
Income from
Operations
Net Cash
Flow
1 $38,000 $119,000 $80,000 $190,000
2 38,000 119,000 61,000 161,000
3 38,000 119,000 30,000 113,000
4 38,000 119,000 13,000 77,000
5 38,000 119,000 6,000 54,000
Total $190,000 $595,000 $190,000 $595,000

Each project requires an investment of $400,000. Straight-line depreciation will be used, and no residual value is expected. The committee has selected a rate of 10% for purposes of the net present value analysis.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1a. Compute the average rate of return for each investment. If required, round your answer to one decimal place.

Average Rate of Return
Front-End Loader %
Greenhouse %

1b. Compute the net present value for each investment. Use the present value of $1 table above. If required, use the minus sign to indicate a negative net present value.

Front-End Loader Greenhouse
Present value of net cash flow $ $
Amount to be invested $ $
Net present value $ $

2. Prepare a brief report for the capital investment committee, advising it on the relative merits of the two investments.

The front-end loader has a   net present value because cash flows occur   in time compared to the greenhouse. Thus, if only one of the two projects can be accepted, the   would be the more attractive.

Feedback

1a. Divide the estimated average annual income by the average investment.

1b. For each investment, multiply the present value factor for each year (Exhibit 2) by that year's net cash flow. Subtract the amount to be invested from the total present value of the net cash flow. Which investment offers the more favorable net present value?

2. Consider when cash flows are received and the time value of money.

Learning Objective 2, Learning Objective 3.
Answer #1

Average Rate of Return = Average Net Income/Average Investment

Average Investment = (Beginning Investment + Ending Investment)/2

= (400,00+0)/2

= $200,000

Average rate of return

Front end Loader = 38,000/200,000

= 19%

Greenhouse = 38,000/200,000

= 19%

1b. NPV = Present value of cash inflows – Present value of cash outflows

Front end Loader

Greenhouse

Present Value of net cash flow

119,000*PVAF(10%, 5 years)

= 119,000*3.79 = $451,010

476,684

Amount to be invested

400,000

400,000

Net Present Value

51,010

76,684

Lower net present value

Later in time

Greenhouse would be more attractive